# Stochastic Population Update Provably Needs An Archive in Evolutionary Multi-objective Optimization

### v1

@@ Abstract
Evolutionary algorithms (EAs) have been widely applied to multi-objective optimization, due to their nature of population-based search. Population update, a key component in multi-objective EAs (MOEAs), is usually performed in a greedy, deterministic manner. However, recent studies have questioned this practice and shown that stochastic population update (SPU), which allows inferior solutions have a chance to be preserved, can help MOEAs jump out of local optima more easily. While introducing randomness in the population update process boosts the exploration of MOEAs, there is a drawback that the population may not always preserve the very best solutions found, thus entailing a large population. Intuitively, a possible solution to this issue is to introduce an archive that stores the best solutions ever found. In this paper, we theoretically show that using an archive can allow a small population and accelerate the search of SPU-based MOEAs substantially. Specifically, we analyze the expected running time of two well-established MOEAs, SMS-EMOA and NSGA-II, with SPU for solving a commonly studied bi-objective problem OneJumpZeroJump, and prove that using an archive can bring (even exponential) speedups. The comparison between SMS-EMOA and NSGA-II also suggests that the $(\mu+\mu)$ update mode may be more suitable for SPU than the $(\mu+1)$ update mode. Furthermore, our derived running time bounds for using SPU alone are significantly tighter than previously known ones. Our theoretical findings are also empirically validated on a well-known practical problem, the multi-objective traveling salesperson problem. We hope this work may provide theoretical support to explore different ideas of designing algorithms in evolutionary multi-objective optimization.
@@ 1 Introduction
Evolutionary algorithms (EAs), a kind of randomized heuristic optimization algorithms inspired by natural evolution, have been found well-suited to MOPs due to their population-based nature. Their widespread applications are across various real-world domains Deb (2001); Coello Coello et al. (2007); Zhou et al. (2019). Notably, there have been developed a multitude of well-established multi-objective EAs (MOEAs), including the non-dominated sorting genetic algorithm II (NSGA-II) Deb et al. (2002), multi-objective evolutionary algorithm based on decomposition (MOEA/D) Zhang and Li (2007), and $\mathcal{S}$ metric selection evolutionary multi-objective optimization algorithm (SMS-EMOA) Beume et al. (2007).
@@
SMS-EMOA $O(\mu n^{k}\cdot\min\{1,\mu/2^{k/4}\})$ Bian et al. (2024b)
@@
[$\mu\geq 2(n-2k+4)$; $p_{s}=1/2$] $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$ [Theorem 3]
$O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$ [Theorem 1] [$\mu\geq 3$; $C=e\mu/(p_{s}(1-p_{c}))$]
[$\mu\geq(n-2k+4)/(1-p_{s})$; $C=e\mu/(p_{s}(1-p_{c}))$]
NSGA-II $O(\mu k(n/2)^{k})$ Bian et al. (2024b)
[$\mu\geq 8(n-2k+3)$; $p_{s}=1/4$; $k>8e^{2}$] $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$ [Theorem 6]
$O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$ [Theorem 4] [$\mu\geq 5$; $C=e/(p_{s}(1-p_{c}))$]
[$\mu\geq 4(n-2k+3)/(1-2p_{s})$; $C=e/(p_{s}(1-p_{c}))$]
Table 1: The expected number of fitness evaluations of SMS-EMOA and NSGA-II for solving OneJumpZeroJump when using SPU alone, or with an archive, where $n$ denotes the problem size, $k$ ($22\leq k<n/$ 2) denotes the parameter of OneJumpZeroJump, $\mu$ denotes the population size, $p_{c}$ denotes the probability of using crossover, and $p_{s}$ denotes the proportion of the current population and offspring solution(s) that SPU selects to preserve directly. The required ranges of $p_{c}$ are: $p_{c}=O(1)$ for Theorems 1 and 4; $p_{c}=\Theta(1)$ for Theorems 3 and 6. The required ranges of $p_{s}$ are: $p_{s}\in[1/(\mu+1),1-1/o(\mu))$ for Theorem 1; $p_{s}\in[1/(\mu+1),(\mu-2)/(\mu+1)]$ for Theorem 3; $p_{s}\in[1/(2\mu),1/2-o(1/\mu))$ for Theorem 4; $p_{s}\in[1/(2\mu),(\mu-4)/(2\mu)]$ for Theorem 6.
@@
Very recently, a study Bian et al. (2023) analytically showed that introducing randomness in the population update process of MOEAs (called stochastic population update or SPU) can help the search. Specifically, the study proved that for SMS-EMOA solving the common benchmark problem OneJumpZeroJump, when $k=n/2-\Omega(n)$, using SPU can bring an acceleration of $\Omega(2^{k/4}/\mu^{2})=\Omega(2^{k/4}/(n-2k+4)^{2})$ on the expected running time, where $n$ denotes the problem size, $k$ ($22\leq k<n/$ 2) denotes the parameter of OneJumpZeroJump, and $\mu$ denotes the population size. Subsequently, Zheng and Doerr Zheng and Doerr (2024b) extended SMS-EMOA to solve a many-objective problem, $m$OneJumpZeroJump, showing that the same SPU can bring an acceleration of $\Theta(2^{k}/\mu)=\Theta(2^{k}/(2n/m-2k+3)^{m/2})$ as well, where $m$ is the number of objectives. These works echoed the empirical studies that show the benefit of considering non-elitism in MOEAs Tanabe and Ishibuchi (2019); Liang et al. (2023a).
SPU in Bian et al. (2023); Zheng and Doerr (2024b) introduces randomness by randomly selecting a proportion $p_{s}$ of the combined set of the current population and offspring solution(s) to be directly preserved into the next generation. This essentially gives a chance for inferior solutions to survive, which enables the evolutionary search to go along inferior regions which may be close to Pareto optimal regions. However, a cost of this method is that there is less space for the best solutions in the population. Some very best solutions to the problem (e.g., globally non-dominated solutions), even found by an MOEA, may be discarded during the population update process. This necessitates a large population used. Unfortunately, when the population size is large, the benefit of using SPU may vanish. This is because the benefit of SPU comes from the operation on inferior solutions while the large population size will lead to a small probability of selecting these solutions. For example, for SMS-EMOA solving OneJumpZeroJump, the acceleration of $\Omega(2^{k/4}/\mu^{2})$ brought by SPU Bian et al. (2023) will vanish when the population size $\mu$ is exponential w.r.t. $k$, e.g., $k=\log n$; for SMS-EMOA solving $m$OneJumpZeroJump, the acceleration of $\Theta(2^{k}/(2n/m-2k+3)^{m/2})$ Zheng and Doerr (2024b) will vanish when the number $m$ of objectives is large, e.g., $m\geq k$, because the population size $\mu=(2n/m-2k+3)^{m/2}$ increases rapidly with $m$. This dilemma impacts both the effectiveness and practicality of using SPU in MOEAs.
@@
  • We theoretically show that incorporating an archive mechanism with SPU can lead to (even exponential) acceleration. For example, comparing the results of Theorems 1 and 3 in Table 1, the expected running time of SMS-EMOA with SPU for solving OneJumpZeroJump, no matter whether an archive is used or not, is $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$, where $C=e\mu/(p_{s}(1-p_{c}))$. The key difference is that using an archive allows for a constant population size. For the case of using SPU alone, when $k=e\ln(8en)$, $p_{s}=1/2$, $p_{c}=1/2$, and $\mu=2n$, the expected running time is $O(n^{k+1})$, which is the same as $O(n^{k+1})$ without SPU Bian et al. (2024b). However, introducing an archive and employing a population size $\mu=5$ reduces the expected running time to $O(n^{k}\cdot(e\ln(20e)/k)^{k-1})$, achieving superpolynomial acceleration and revitalizing the effectiveness of SPU. Even when using SPU alone is effective (e.g., the expected running time is $O(n^{k+1}(8e\ln(8en)/n)^{n/8-1})$ when $k=n/8$, $p_{s}=1/2$, $p_{c}=1/2$, and $\mu=2n$), using an archive can further bring exponential acceleration (e.g., the expected running time reduces to $O(n^{k}(8e\ln(20e)/n)^{n/8-1})$ by setting $\mu=5$). Note that Bian et al. Bian et al. (2024a) recently proved the effectiveness of using an archive (bringing polynomial acceleration) for MOEAs, while our analysis for MOEAs with SPU reveals that even exponential acceleration can be obtained.
@@
  • Comparing the results of Theorems 1 and 4 in Table 1, we can find that the expected running time for both SMS-EMOA and NSGA-II to solve OneJumpZeroJump is $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$. The key difference lies in that $C=e\mu/(p_{s}(1-p_{c}))$ for SMS-EMOA, whereas $C=e/(p_{s}(1-p_{c}))$ for NSGA-II. This implies that NSGA-II is better since a smaller $C$ leads to a shorter running time. Our analysis reveals that the benefit of NSGA-II is due to its $(\mu+\mu)$ update mode, which selects each solution in the current population for reproduction and thus makes exploring promising dominated solutions easier. Similar results can be observed by comparing Theorems 3 and 6. Thus, we suggest that the $(\mu+\mu)$ update mode may be more suitable for SPU than the $(\mu+1)$ update mode.
  • In addition, our derived running time bounds for MOEAs with SPU in the second column of Table 1 significantly improve the previously known ones Bian et al. (2024b); Zheng and Doerr (2024b). For NSGA-II solving OneJumpZeroJump, when using SPU with $p_{s}=1/4$ and $p_{c}=0$, the expected running time derived in Bian et al. (2024b) is $O(\mu k(n/2)^{k})$, whereas our result in Theorem 4 is $O(\mu n^{k}\cdot(e\ln(4e)/k)^{k-1})$ with the same settings. This implies an improvement ratio of $\Theta(k(k/(2e\ln(4e)))^{k-1})$, which can be exponential when $k$ is large, e.g., $k=n/8$. Similar improvements can be observed for SMS-EMOA with SPU by comparing the result in Theorem 1 with those in Bian et al. (2024b); Zheng and Doerr (2024b). Moreover, our running time bounds are more general, as we consider variable survival probability $p_{s}$ and crossover probability $p_{c}$.
@@
We also validate our theoretical findings through an empirical study on a practical problem, the multi-objective travelling salesperson problem [Ribeiro et al., 2002]. The results show that combing SPU with archive leads to the best performance for both SMS-EMOA and NSGA-II. Furthermore, with the SPU method, NSGA-II always performs better than SMS-EMOA. These results confirm our theoretical findings.
@@
Finally, we give a brief overview about the running time analysis of MOEAs. Over the last decade, there has been an increasing interest for the evolutionary theory community to study MOEAs. Primitive theoretical works Giel (2003); Laumanns et al. (2004a, b); Neumann (2007); Giel and Lehre (2010); Neumann and Theile (2010); Doerr et al. (2013); Qian et al. (2013, 2016); Bian et al. (2018) mainly focus on analyzing the expected running time of a simple MOEA like GSEMO/SEMO. Recently, researchers have begun to examine practical MOEAs. Huang et al. Huang et al. (2021) investigated MOEA/D, assessing the effectiveness of different decomposition methods. Zheng et al. Zheng et al. (2022) conducted the first analysis of the expected running time of NSGA-II. Bian et al. Bian et al. (2023) analyzed the running time of SMS-EMOA and showed that SPU can bring acceleration. Moreover, Wietheger and Doerr Wietheger and Doerr (2023) demonstrated that NSGA-III Deb and Jain (2014) exhibits superior performance over NSGA-II in solving the tri-objective problem $3$OneMinMax. Ren et al. Ren et al. (2024a) analyzed the running time of SPEA2 on three $m$-objective problems. Some other works on well-established MOEAs include Bian and Qian (2022); Zheng and Doerr (2022, 2024a); Cerf et al. (2023); Dang et al. (2023a, b); Doerr and Qu (2023b, c); Doerr et al. (2024a, b); Lu et al. (2024); Opris et al. (2024); Opris (2024); Ren et al. (2024b).
@@ 2 Multi-objective Optimization
Next, we introduce the benchmark problem OneJumpZeroJump studied in this paper. The OneJumpZeroJump problem as presented in Definition 3 is constructed based on the Jump problem Doerr and Neumann (2020), and has been widely used in MOEAs’ theoretical analyses Doerr and Zheng (2021); Doerr and Qu (2023a); Lu et al. (2024); Ren et al. (2024b). Its first objective is the same as the Jump problem, which aims to maximize the number of 1-bits of a solution, except for a valley around $1^{n}$ (the solution with all 1-bits) where the number of 1-bits should be minimized. The second objective is isomorphic to the first, with the roles of 1-bits and 0-bits reversed. The Pareto front of the OneJumpZeroJump problem is $\{(a,n+2k-a)\mid a\in[2k..n]\cup\{k,n+k\}\}$, whose size is $n-2k+3$, and the Pareto optimal solution corresponding to $(a,n+2k-a)$, $a\in[2k..n]\cup\{k,n+k\}$, is any solution with $(a-k)$ 1-bits. We use $F_{I}^{*}=\{(a,n+2k-a)\mid a\in[2k..n]\}$ to denote the inner part of the Pareto front.
@@ 3 Running Time Analysis of SMS-EMOA
In this section, we prove that using an archive can bring (even exponential) speedup for the well-established MOEA, SMS-EMOA Beume et al. (2007), with stochastic population update (SPU) solving the OneJumpZeroJump problem. In Section 3.1, we first introduce SMS-EMOA with SPU, and its expected running time for solving OneJumpZeroJump, which is much tighter than the previous results Bian et al. (2024b); Zheng and Doerr (2024b). Then, in Section 3.2, we prove that using an archive can significantly reduce the running time.

### v2

# A Theoretical Perspective on Why Stochastic Population Update Needs an Archive in Evolutionary Multi-objective Optimization

@@ Abstract
Evolutionary algorithms (EAs) have been widely applied to multi-objective optimization due to their population-based nature. Population update, a key component in multi-objective EAs (MOEAs), is usually performed in a greedy, deterministic manner. However, recent studies have questioned this practice and shown that stochastic population update (SPU), which allows inferior solutions have a chance to be preserved, can help MOEAs jump out of local optima more easily. Nevertheless, SPU risks losing high-quality solutions, potentially requiring a large population. Intuitively, a possible solution to this issue is to introduce an archive that stores the best solutions ever found. In this paper, we theoretically show that using an archive allows a small population and may enhance the search performance of SPU-based MOEAs. We examine two classic algorithms, SMS-EMOA and NSGA-II, on the bi-objective problem OneJumpZeroJump, and prove that using an archive can reduce the expected running time upper bound (even exponentially). The comparison between SMS-EMOA and NSGA-II also suggests that the $(\mu+\mu)$ update mode may be more suitable for SPU than the $(\mu+1)$ update mode. We also validate our findings empirically. We hope this work may provide theoretical support to explore different ideas of designing algorithms in evolutionary multi-objective optimization.
@@ 1 Introduction
Evolutionary algorithms (EAs), a kind of randomized heuristic optimization algorithms inspired by natural evolution, have been found well-suited to MOPs due to their population-based nature. Their widespread applications are across various real-world domains Deb (2001); Zhou et al. (2019). Notably, there have been developed a multitude of well-established multi-objective EAs (MOEAs), including the non-dominated sorting genetic algorithm II (NSGA-II) Deb et al. (2002), multi-objective evolutionary algorithm based on decomposition (MOEA/D) Zhang and Li (2007), and $\mathcal{S}$ metric selection evolutionary multi-objective optimization algorithm (SMS-EMOA) Beume et al. (2007).
@@
SMS-EMOA $O(\mu n^{k}\cdot\min\{1,\sqrt{k}\mu/2^{k/2}\})$ Bian et al. (2024b)
@@
$\mu\geq 2(n-2k+4)$; $p_{s}=1/2$] $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$ [Theorem 3]
$O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$ [Theorem 1] $\mu\geq 3$; $C=e\mu/(p_{s}(1-p_{c}))$]
$\mu\geq(n-2k+4)/(1-p_{s})$; $C=e\mu/(p_{s}(1-p_{c}))$]
NSGA-II $O(\mu\sqrt{k}(n/2)^{k})$ Bian et al. (2024b)
$\mu\geq 8(n-2k+3)$; $p_{s}=1/4$; $k>8e^{2}$] $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$ [Theorem 6]
$O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$ [Theorem 4] $\mu\geq 5$; $C=e/(p_{s}(1-p_{c}))$]
$\mu\geq 4(n-2k+3)/(1-2p_{s})$; $C=e/(p_{s}(1-p_{c}))$]
Table 1: The expected number of fitness evaluations of SMS-EMOA and NSGA-II for solving OneJumpZeroJump when using SPU alone, or with an archive, where $n$ denotes the problem size, $k$ ($2\leq k<n/2$) denotes the parameter of OneJumpZeroJump, $\mu$ denotes the population size, $p_{c}$ denotes the probability of using crossover, and $p_{s}$ denotes the proportion of the current population and offspring solution(s) that SPU selects to preserve directly. The required ranges of $p_{c}$ are: $1-p_{c}=\Omega(1)$ for Theorems 1 and 4; $p_{c}=\Theta(1)$ for Theorems 3 and 6. The required ranges of $p_{s}$ are: $p_{s}\in[1/(\mu+1),1-o(1/\mu))$ for Theorem 1; $p_{s}\in[1/(\mu+1),(\mu-2)/(\mu+1)]$ for Theorem 3; $p_{s}\in[1/(2\mu),1/2-o(1/\mu))$ for Theorem 4; $p_{s}\in[1/(2\mu),(\mu-4)/(2\mu)]$ for Theorem 6.
@@
Very recently, Bian et al. Bian et al. (2024b) analytically showed that introducing randomness in the population update process of MOEAs (called stochastic population update, SPU) can help the search. Specifically, the study proved that for SMS-EMOA solving the common benchmark problem OneJumpZeroJump, when $k=n/2-\Omega(n)$, using SPU can bring an acceleration of $\Omega(2^{k/2}/(\sqrt{k}\mu^{2}))=\Omega(2^{k/2}/(\sqrt{k}(n-2k+4)^{2}))$ on the expected running time, where $n$ denotes the problem size, $k$ ($2\leq k<n/2$) denotes the parameter of OneJumpZeroJump, and $\mu$ denotes the population size. Subsequently, Zheng and Doerr Zheng and Doerr (2024b) extended SMS-EMOA to solve a many-objective problem, mmOneJumpZeroJump, showing that the same SPU can bring an acceleration of $\Theta(2^{k}/\mu)=\Theta(2^{k}/(2n/m-2k+3)^{m/2})$ as well, where $m$ is the number of objectives. These works echoed the empirical studies that show the benefit of considering non-elitism in MOEAs Tanabe and Ishibuchi (2019); Liang et al. (2023a).
SPU in Bian et al. (2024b); Zheng and Doerr (2024b) introduces randomness by randomly selecting a proportion $p_{s}$ of the combined set of the current population and offspring solution(s) to be directly preserved into the next generation. This essentially gives a chance for inferior solutions to survive, which enables the evolutionary search to go along inferior regions which may be close to Pareto optimal regions. However, a cost of this method is that there is less space for the best solutions in the population. Some very best solutions to the problem (e.g., globally non-dominated solutions), even found by an MOEA, may be discarded during the population update process. This necessitates a large population used. Unfortunately, when the population size is large, the benefit of using SPU may vanish. This is because the benefit of SPU comes from the operation on inferior solutions while the large population size will lead to a small probability of selecting these solutions. For example, for SMS-EMOA solving OneJumpZeroJump, the acceleration of $\Omega(2^{k/2}/(\sqrt{k}\mu^{2}))$ brought by SPU Bian et al. (2024b) will vanish when the population size $\mu$ is exponential w.r.t. $k$, e.g., $k=\log n$; for SMS-EMOA solving mmOneJumpZeroJump, the acceleration of $\Theta(2^{k}/(2n/m-2k+3)^{m/2})$ Zheng and Doerr (2024b) will vanish when the number $m$ of objectives is large, e.g., $m\geq k$, because the population size $\mu=(2n/m-2k+3)^{m/2}$ increases rapidly with $m$. This dilemma impacts both the effectiveness and practicality of using SPU in MOEAs.
@@
  • We theoretically show that incorporating an archive mechanism with SPU can reduce the upper bound on the expected running time (even exponentially). For example, comparing Theorems 1 and 3 in Table 1, the expected running time of SMS-EMOA with SPU for solving OneJumpZeroJump, no matter whether an archive is used or not, is $O(\mu n^{k}\cdot\min\{1,(e\ln C/k)^{k-1}\})$, where $C=e\mu/(p_{s}(1-p_{c}))$. The key difference is that using an archive allows a constant population size, resulting in a significantly smaller $C$ and thus reducing the upper bound significantly. Note that Bian et al. Bian et al. (2024a) recently proved the effectiveness of using an archive (bringing polynomial acceleration) for MOEAs, while our analysis for MOEAs with SPU reveals that even exponential acceleration can be obtained.
@@
  • Comparing the results of Theorems 1 and 4 in Table 1, we can find that the upper bound of NSGA-II is smaller than SMS-EMOA when using SPU. Our analysis reveals that the benefit of NSGA-II is due to its $(\mu+\mu)$ update mode, which selects each solution in the current population for reproduction and thus makes exploring promising dominated solutions easier.
  • In addition, our derived running time bounds for MOEAs with SPU in the second column of Table 1 are significantly better than the previously known ones Bian et al. (2024b); Zheng and Doerr (2024b). This improvement stems from the analysis method of considering both the number and size of jumps across gap of dominated solutions. The method happens to share similarities with that of proving Lemma 12 in doerr2024hyper. Moreover, our running time bounds are more general, as we consider variable survival probability $p_{s}$ and crossover probability $p_{c}$.
@@
We also validate our theoretical findings through an empirical study on the artificial OneJumpZeroJump problem and the multi-objective travelling salesperson problem (MOTSP) Ribeiro et al. (2002). The results show that combining SPU with an archive leads to the best performance for both SMS-EMOA and NSGA-II. Furthermore, the results on the MOTSP show that with the SPU method, NSGA-II always performs better than SMS-EMOA. These results confirm our theoretical findings.
@@
Finally, we give a brief overview about the running time analysis of MOEAs. Over the last decade, there has been an increasing interest for the evolutionary theory community to study MOEAs. Early theoretical works Laumanns et al. (2004a, b); Neumann (2007); Giel and Lehre (2010); Neumann and Theile (2010); Doerr et al. (2013); Qian et al. (2013, 2016); Bian et al. (2018) mainly focus on analyzing the expected running time of a simple MOEA like GSEMO/SEMO. Recently, researchers have begun to examine practical MOEAs. Huang et al. Huang et al. (2021) investigated MOEA/D, assessing the effectiveness of different decomposition methods. Zheng et al. Zheng et al. (2022) conducted the first theoretical analysis of NSGA-II. Bian et al. Bian et al. (2024b) analyzed the running time of SMS-EMOA and showed that SPU can bring acceleration. Moreover, Wietheger and Doerr Wietheger and Doerr (2023) demonstrated that NSGA-III Deb and Jain (2014) exhibits superior performance over NSGA-II in solving the tri-objective problem 33OneMinMax. Ren et al. Ren et al. (2024a) analyzed the running time of SPEA2 on three $m$-objective problems. Some other works on well-established MOEAs include Bian and Qian (2022); Zheng and Doerr (2022, 2024a); Cerf et al. (2023); Dang et al. (2023a, b); Doerr and Qu (2023b, c); Doerr et al. (2024a, b); Lu et al. (2024); Opris et al. (2024); Opris (2024); Ren et al. (2024b).
@@ 2 Multi-objective Optimization
Note that the notions of “weakly dominate” and “dominate” are also called “dominate” and “strictly dominate” in some works Cerf et al. (2023); Wietheger and Doerr (2023).
Next, we introduce the benchmark problem OneJumpZeroJump studied in this paper. The OneJumpZeroJump problem as presented in Definition 3 is constructed based on the Jump problem Doerr and Neumann (2020), and has been widely used in MOEAs’ theoretical analyses Doerr and Zheng (2021); Doerr and Qu (2023a); Lu et al. (2024); Ren et al. (2024b). Its first objective is the same as the Jump problem, which aims to maximize the number of 1-bits of a solution, except for a valley around $1^{n}$ (the solution with all 1-bits) where the number of 1-bits should be minimized. The second objective is isomorphic to the first, with the roles of 1-bits and 0-bits reversed. The Pareto front of the OneJumpZeroJump problem is $\{(a,n+2k-a)\mid a\in[2k..n]\cup\{k,n+k\}\}$, whose size is n2k+3n-2k+3, and the Pareto optimal solution corresponding to (a,n+2ka)(a,n+2k-a), $a\in[2k..n]\cup\{k,n+k\}$, is any solution with (ak)(a-k) 1-bits. We use $F_{I}^{*}=\{(a,n+2k-a)\mid a\in[2k..n]\}$ to denote the inner part of the Pareto front.
@@ 3 Running Time Analysis of SMS-EMOA
In this section, we prove that using an archive can bring (even exponential) speedup for the well-established MOEA, SMS-EMOA Beume et al. (2007), with stochastic population update (SPU) solving the OneJumpZeroJump problem. In Section 3.1, we first introduce SMS-EMOA with SPU, and its expected running time for solving OneJumpZeroJump, which is much tighter than the previous results Bian et al. (2024b); Zheng and Doerr (2024b). Then, in Section 3.2, we prove that using an archive can significantly reduce the upper bound on the expected running time. Due to space limitation, all mathematical proofs could only be sketched or had to be omitted completely.